CH13 Q8

A two-way analysis of variance experiment with interaction was conducted. Factor A had three levels (columns), factor B had five levels (rows), and six observations were obtained for each combination. The results include the following sum of squares terms: SST 1548 SSA 1022 SSB 390 SSAB 26 a. Construct an ANOVA table. (Round intermediate calculations to at least 4 decimal places. Round MS to 4 decimal places and F to 3 decimal places.) Answer is not complete. ANOVA Source of Variation Rows Columns Interaction Error Total p- 0.000 0.000 0.035 df MS value
Answer #1

Solution :-

Given data :

SST = 1548

SSA = 1022

SSB = 390

SSAB = 26

Factor A had three levels (columns), factor B had five levels (rows), and six observations were obtained for each combination.

Here we need to construct the ANOVA table.

ANOVA
Source of variation SS DF MS F P-Value
Rows ( B ) 390

=(5 - 1)

= 4

= (395 / 4)

= 98.75

= (98.75 /1.467)

= 67.314

0.000
Columns ( A ) 1022

=(3 - 1)

= 2

= (1022 / 2)

= 511

= (511 / 1.467)

=348.321

0.000
Interaction ( A * B ) 26

=(4 * 2)

= 8

= (26 / 8)

= 3.25

= (3.25 / 1.467)

= 2.215

0.035
Error

=(1548-(390+1022+26))

=110

=(5 * 3 * (6-1))

= (5 * 3 * 5)

= 75

= (110 / 75)

=1.467

Total 1548 89